The total length of 3 pipes is 24 inches. The middle-sized pipe is 6 inches longer than the smallest. The largest is twice as long as the smallest. How long is the middle-sized pipe?
Answer by ikleyn(52781)

.

The middle   size = x

the smallest size = x-6

the longest  size = 2(x-6)


The total length equation is as simple as


    x + (x-6) + 2(x-6) = 24   inches.


You do the rest.

Solved.